OpenStudy (anonymous):

find the shaded area. When appropriate, use the pie key on your calculator and round your answer to the nearest hundredth. 5in 10 in-shaded area (((find the area of the bigger semi-circle, subtract the area of the smaller semi-circle from it to get the area of the shaded region))

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

smh

OpenStudy (anonymous):

\[ A = \pi(10^{2} - 5^{2}) = 75\pi \approx 235.62 \] square inches.

OpenStudy (anonymous):

how do i make the symbols?

OpenStudy (anonymous):

and wats the formula pls

OpenStudy (anonymous):

so a semicircle is ( π r^2 ) /2 assumption #1 radius of large circle is 10, smaller is 5 so shaded area = (πR^2 -πr^2)/2 = π(10^2 -5^2) /2 = 235.619 /2 = 117.8097.